OpenStudy (anonymous):

Convert r= 1/ (1+SinƟ) into rectangular coordinates.

OpenStudy (jdoe0001):

rectangular coordinates.<--- or expression?

OpenStudy (anonymous):

Directions in the book say "Convert the polar equation to rectangular coordinates."

OpenStudy (jdoe0001):

well... we don't have an angle to use though

OpenStudy (anonymous):

I figured that you could multiply both sides by r, which would become r^2= r/ (r+rsinƟ). rsinƟ= y

OpenStudy (jdoe0001):

yes, but that'd give you a "rectangular expression" or "equation", no a coordinate set

OpenStudy (anonymous):

I don't need actual numbers. I just have to simplify the equation into terms of x and y

OpenStudy (jdoe0001):

so \(\bf r=\cfrac{1}{1+sin(\theta)}\implies r[1+sin(\theta)]=1\implies r+rsin(\theta)=1\\ \quad \\ {\color{blue}{ r^2=x^2+y^2\implies r=\sqrt{x^2+y^2}\qquad rsin(\theta)=y}}\\ \quad \\ r+rsin(\theta)=1\implies \sqrt{x^2+y^2}+y=1\implies \sqrt{x^2+y^2}=1-y\\ \quad \\ (\sqrt{x^2+y^2})^2=(1-y)^2\)

OpenStudy (jdoe0001):

so.... you'd just expand the binomial on the right-side, and simplify

OpenStudy (dumbcow):

\[r + r \sin \theta = 1\] \[r+y = 1\] \[r^2 = (1-y)^2\] \[x^2 +y^2 = (1-y)^2\]

OpenStudy (anonymous):

jdoe, I understand that you multiplied (1+sinƟ) by r,but how did you get that = to 1?

OpenStudy (jdoe0001):

hmm the 1 was given

OpenStudy (jdoe0001):

\(\bf r=\cfrac{{\color{red}{ 1}}}{1+sin(\theta)}\implies r[1+sin(\theta)]={\color{red}{ 1}}\implies r+rsin(\theta)={\color{red}{ 1}}\)

OpenStudy (anonymous):

You cross-multiplied. Woops, my bad.